For the floor plans given in exercise 27, determine whether the side through the points (2,3) and (11,6) is perpendicular to the side through the points (2,3) and (-3,18). The floor plans for the building (in exercise 27) have the four corners of a room located at the points (2,3), (11,6), (-3,18), and (8,21).
McGraw Hill 6th edition, 7.2, Problem # 28. Thanks!!

Answer by stanbon(75887)

determine whether the side through the points (2,3) and (11,6) is perpendicular to the side through the points (2,3) and (-3,18).
-------------
Slope #1 = [6-3]/[11-2]= 3/9=1/3
Slope #2 = [18-3]/[-3-2]=15/-5 =-3
Since each is the negative reciprocal of the other the lines are
perpendicular.
Cheers,
Stan H.